Find the slope of the line that passes through the points (-9, -9) and (4, -1)
Answer by checkley77(12844)

(-9, -9) and (4, -1)
SLOPE=(Y2-Y1)/(x2-X1)=(-1+9)/(4+9)=8/13 ANS. FOR THE SLOPE.